I need 3 simple examples of non factorable quadratic equations. (so you have to use the quadratic formula)
Answer by longjonsilver(2297)

x^2 -5x + 3 = 0
2x^2 -6x - 2 = 0
x^2 +3x - 1 = 0
jon